# Real-time Localization and Mapping in Architectural Plans with Deviations

### v1

@@ VI Graph Optimization
When the robot starts navigating the environment having a prior A-Graph, it also generates in real-time an S-Graph. The overall state at time can be defined as:
where are the robot poses at selected keyframes in the S-Graph frame of reference, are the plane parameters of the and wall-surfaces of the S-Graph and A-Graph respectively, contains the parameters of the and four-wall rooms of the S-Graph and A-Graph respectively. are the parameters of the two-wall rooms in the S-Graph. contains the parameters of doorways of the A-Graph, are the floors levels, and models the drift between the odometry frame and the S-Graph reference frame . If at time there is no match obtained between the A-Graph and S-Graph we perform single S-Graph optimization as explained in [4].
@@
Transformation Estimation: Upon receiving the match (Section. IV) and performing graph merging (Section. V), we augment our global state with additional transformation factor . represents the transformation between the origins of the A-Graph and the S-Graph and add a factor between the two origins of the graph. It is important to note that at this stage, the wall-surface and room entities with possible deviation are not included for estimating .
@@
Combined Optimization. Finally, after getting the initial estimates of the transformation between the origins and the deviations between the semantic entities, we optimize the whole state to simultaneously estimate the position of each semantic entity, deviations, and the transformation between the two graphs.
@@ VII-A Methodology
Baselines. We have selected the following LiDAR-based baselines for comparisons due to their suitability, their reported results, and the availability of their code: AMCL [6], UKFL [10], OGM2PGM [9], IR-MCL [11], and iS-Graphs [5]. As each baseline takes a different map input for localization, we used 2D occupancy grid maps for AMCL and OGM2PGM, 3D meshes for UKFL and IR-MCL, and A-Graphs for iS-Graphs and our method diS-Graphs, respectively.
Simulated Datasets. We validate the algorithms in five simulated datasets named SE1 to SE5. To record the datasets, we use Gazebo physics simulator to recreate the robot, its sensors (LiDAR), and the 3D indoor environments obtained from actual architectural plans. We report the absolute trajectory error (ATE) compared with the available ground truth trajectory. We also report the localization convergence success rate of all methods.
Real Datasets. We collected data with a legged robot equipped with a Velodyne VLP-16 LiDAR, at five different construction sites (RE1 to RE5), with existing architectural plans. In real experiments, we report the Root Mean Square Error (RMSE) of the estimated 3D maps against the non-deviated 3D maps from the architectural plan. Furthermore, we report all methods’ convergence rates and convergence and computation times.
Deviations. Note that, given the absence of actual ground truth deviations in a real environment with respect to the plans, we explicitly deviate the wall-surfaces entities in the architectural plans to have a ground truth estimate of the deviations for both simulated and real datasets. For all datasets, we performed five separate tests introducing uniform random wall-surface deviations in architectural plans ranging from to in translation, and and in rotations.
@@ VII-B Results and Discussion
Absolute Trajectory Error. Table I shows the ATE for all baselines and our diS-Graphs. Our method outperforms all baselines. Specifically, it shows an error reduction of around compared to AMCL, compared to OGM2PGM, compared to IR-MCL, compared to UKFL, and a reduction of over iS-Graphs, which uses architectural plans but assumes no deviations.
@@
Fig. 6(a) summarizes the ATE performance of all the baseline algorithms. AMCL shows the highest median ATE and a relatively narrow distribution, indicating consistently high error rates. UKFL and OGM2PGM demonstrate moderate performance with similar median ATEs, although OGM2PGM shows a wider range of errors. IR-MCL exhibits the largest variability, suggesting inconsistent performance in different scenarios. Our diS-Graphs shows the lowest median ATE and the most compact distribution, indicating consistently low error rates under various conditions. This shows that the addition of explicit deviation factors between semantic elements reduces not only the ATE but also its variance.
Pointcloud Alignment Error. Table II shows the RMSE of the point clouds with respect to the ground truth for all methods and ours. In case of deviations in construction from the plans, diS-Graphs shows better accuracy than AMCL, better than OGM2PGM, better than IR-MCL, and better than iS-Graphs. Although UKFL’s average error is equal to diS-Graphs’, it has a very low convergence rate (see Table III) rendering the comparison unfair. Moreover, it cannot estimate the deviations between “as-planned” and “as-built” environments. Fig. 6(b) summarizes the performance of all algorithms in real environments. AMCL exhibits the highest median and widest interquartile range, indicating greater variability in performance. IR-MCL presents a large spread of results, while OGM2PGM shows moderate performance with a smaller range of variability compared to AMCL and IR-MCL. UKFL and diS-Graphs show the lowest median RMSE, suggesting superior accuracy. Because of our simultaneous estimation of deviations and initial transformation, we can not only simultaneously globally localize the robot but also estimate the deviations between semantic elements of “as-planned” and “as-built” environments moreover improving the overall mapping accuracy compared to other algorithms.
@@
Convergence Rate. In terms of localization convergence rate in simulated environments (Table III) our approach shows a convergence rate better than AMCL, better than UKFL and better than OGM2PGM. Although the convergence rate of our algorithm is on par with that of IR-MCL, IR-MCL shows inconsistent performance in terms of ATE. We observed that since AMCL and UKFL rely heavily on the initial estimate of the robot’s starting position, their convergence rate is affected in environments with complex geometries, whereas the accuracy of OGM2PGM degrades in symmetric environments. IR-MCL, on the other hand, converges on most datasets but requires the model to be trained on every new dataset.
Table III shows the localization convergence rate in real environments. Our method shows the best convergence rate of all algorithms in four real environments (RE1-RE4) because of its ability to detect deviated elements in the environment. Our algorithm fails only in RE5 given the fact that the algorithm could not detect enough rooms in the environment to perform the graph matching.
@@
Convergence and Computation Time. Table IV shows the convergence and computation time for each algorithm. The computation time is the time used for each pose update. On average, the computation time of our algorithm is the best compared to other 3D algorithms, showing its real-time performance. OGM2PGM and AMCL have the best computation time because, unlike others, they process 2D information. The convergence time is the time it takes the algorithm to globally localize. IR-MCL and OGM2PGM have the best convergence time. iS-Graphs and diS-Graphs have considerably longer convergence times because they need to detect a certain number of rooms in the environment for graph matching to find a unique match. In addition, the previous version of graph matching struggled to resolve symmetries during the matching process, resulting in longer convergence times. However, the modifications in the graph matching algorithm (Section IV) proposed in this work improve the symmetry resolution ability and reduce the convergence time by almost .
@@
Deviation Estimation. Fig. 7 shows the amount of deviation our algorithm can correctly estimate in real environments. The accuracy level of our LiDAR is [24]. Considering that of the data falls within of the mean, and the fact that we are trying to estimate the deviations as well as transformation between the two graphs simultaneously, we can only accurately estimate the deviations greater than . The maximum translational and rotational deviation in wall-surfaces our algorithm can detect accurately is and respectively. Fig. 5 shows several qualitative results from our diS-Graphs method for real experiments. The robot can successfully localize, map, and estimate deviations in these environments. Note that we only show the rooms and walls for better understanding, and all the other semantic elements and the robot are not shown. We do not show RE5 in Fig. 5, as the robot could not localize itself in this sequence.
@@
Ablation Study. Table I shows that associating ‘uniform covariance’ in simulated datasets (diS-Graphs (UC)) the algorithm cannot differentiate between deviated and non-deviated elements which results in poor pose estimation. In some cases, the use of uniform covariances results in even worse performance than iS-Graphs. Similarly, when using single optimization (diS-Graphs (SO)) instead of alternating optimization, the algorithm cannot differentiate between the transformation between two graphs and the deviations between their elements, resulting in higher ATE as shown in Table I.
Table II shows the ablation of uniform covariances and single optimization in real-world datasets. Using uniform covariances (diS-Graphs (UC)) and single optimization (diS-Graphs (SO)) results in lower mapping accuracy due to the algorithm’s inability to accurately map the deviated elements and differentiate between initial transformation and deviations simultaneously.
@@ VIII Conclusions and Future Work
In this paper, we present our work on real-time localization and mapping in architectural plans with deviations while simultaneously estimating deviations between structural elements (specifically walls and rooms) of “as-planned” and “as-built” environments. Our work demonstrates higher localization accuracy in simulated environments, and higher mapping accuracy in real environments, and more robust performance in the presence of deviations between “as-planned” and “as-built” environments, compared to the best-performing existing method in the literature. Additionally, our algorithm provides an estimate of existing deviations up to in translation and in rotation.
Our algorithm is limited by the need to have enough distinctive semantic elements (i.e. wall-surfaces and rooms) to provide a unique match. As future work, we plan to add the ability to detect and match more semantic elements, which will translate into an improvement in the convergence rate and deviation detection range. Moreover, to gain flexibility, we plan to improve the matching process by adding the ability to detect rooms consisting of more than four walls.

### v2

@@ VI Graph Optimization
Before the S-Graph and A-Graph are coupled, the S-Graph state at time can be defined as:
Similarly, the A-Graph state is defined as:
where are the robot poses at selected keyframes in the S-Graph frame of reference, are the plane coefficients of the and wall-surfaces of the S-Graph and A-Graph respectively, contains the poses of the and four-wall rooms of the S-Graph and A-Graph respectively. are the poses of the two-wall rooms in the S-Graph. contains the poses of doorways of the A-Graph, are the floors levels, and models the drift between the odometry frame and the S-Graph reference frame . If at time there is no match obtained between the A-Graph and S-Graph we perform single S-Graph optimization as detailed in [7].
@@
Transformation Estimation: Upon receiving the match (Section IV) and performing graph coupling (Section V), we combine Eq. 4 and Eq. 5 to make a global state with additional transformation factor . represents the transformation between the origins of the A-Graph and the S-Graph. It is important to note that at this stage, the deviated wall-surface and room are not included for estimating .
@@
Joint Optimization. Finally, after getting the initial estimates of the transformation between the origins and the deviations between the semantic entities, we optimize the whole state to simultaneously estimate the position of each semantic entity, deviations, and the transformation between the two graphs.
@@ VII-A Methodology
Baselines. We have selected the following LiDAR-based baselines for comparisons due to their suitability, their reported results, and the availability of their code: AMCL [9], UKFL [13], OGM2PGM [12], IR-MCL [14], and iS-Graphs [8]. Although some works do localization in imprecise architectural plans, the absence of open-source implementations prevents direct comparison with these methods. As each selected baseline takes a different map input for localization, we used 2D occupancy grid maps for AMCL and OGM2PGM, 3D meshes for UKFL and IR-MCL, and A-Graphs for iS-Graphs and our method diS-Graphs, respectively.
Simulated Datasets. We validate the algorithms in five simulated datasets named SE1 to SE5. To record the datasets, we use the Gazebo physics simulator to recreate the robot, its sensors (LiDAR), and the 3D indoor environments obtained from actual architectural plans. We report the absolute trajectory error (ATE) compared with the available ground truth trajectory. We also report the localization convergence success rate of all methods.
Real Datasets. We collected data with a legged robot equipped with a Velodyne VLP-16 LiDAR, at five different construction sites (RE1 to RE5), with existing architectural plans. In real experiments, we report the Root Mean Square Error (RMSE) between estimated 3D maps and ground truth architectural plans due to the absence of pose ground truth. The RMSE, while a mapping metric, primarily demonstrates the localization accuracy of all baselines, as mapping accuracy inherently reflects pose estimation quality. Furthermore, we report all methods’ convergence rates and convergence and computation times.
Deviations. Note that, given the absence of actual ground truth deviations in a real environment with respect to the plans, we explicitly deviate the wall-surfaces entities in the architectural plans to have a ground truth estimate of the deviations for both simulated and real datasets. For all datasets, we performed five separate tests introducing uniform random wall-surface deviations in architectural plans ranging from to in translation, and and in rotations. Moreover, to test the robustness of our algorithm against deviations, we conducted 25 experiments per dataset with varying deviations.
@@ VII-B Results and Discussion
Absolute Trajectory Error. Table I shows the average ATE for all baselines and our diS-Graphs in the presence of deviations between “as-planned” and “as-built” environments. Our method shows an error reduction of around compared to AMCL, compared to OGM2PGM, compared to IR-MCL, compared to UKFL, and a reduction of over iS-Graphs.
Bold values are the best and the second best are underlined. ‘-’ refers to an unsuccessful run.
@@
Fig. 6(a) summarizes the ATE performance of all the baseline algorithms. AMCL shows the highest median ATE and a relatively narrow distribution, indicating consistently high error rates. UKFL and OGM2PGM demonstrate moderate performance with similar median ATEs, although OGM2PGM shows a wider range of errors. IR-MCL exhibits the largest variability, suggesting inconsistent performance in different scenarios. AMCL, OGM2PGM, IR-MCL, UKFL and iS-Graphs assume no deviations between “as-planned” and “as-built” environments. These methods attempt to match laser scans with an incorrect reference map, leading to consistent misalignment. Therefore, they show higher ATEs across datasets, indicating their vulnerability to architectural deviations. diS-Graphs can detect and estimate the deviations between “as-planned” and “as-built” environments, resulting in lower ATE than other baselines.
Point Cloud Alignment Error. Table II shows the RMSE of the point clouds with respect to the ground truth for all methods and ours. In case of deviations in construction from the plans, diS-Graphs shows better accuracy than AMCL, better than OGM2PGM, better than IR-MCL, and better than iS-Graphs. Although UKFL’s average error is equal to diS-Graphs’, it has a very low convergence rate (see Table III) rendering the comparison unfair. Moreover, it cannot estimate the deviations between “as-planned” and “as-built” environments. Fig. 6(b) summarizes the performance of all algorithms in real environments. AMCL exhibits the highest median and widest interquartile range, indicating greater variability in performance. IR-MCL presents a large spread of results, while OGM2PGM shows moderate performance with a smaller range of variability compared to AMCL and IR-MCL. Although UKFL and diS-Graphs show the lowest median RMSE, suggesting superior accuracy, the comparison with UKFL is not fair because of its low convergence rate. Because of our simultaneous estimation of deviations and initial transformation, we can not only simultaneously globally localize the robot but also estimate the deviations between semantic elements of “as-planned” and “as-built” environments moreover improving the overall accuracy compared to other algorithms.
@@
Convergence Rate. Our approach demonstrates superior convergence rates across both simulated and real environments (Table III). In simulations, our method achieves a convergence rate better than AMCL, better than UKFL, and better than OGM2PGM. While our convergence rate matches that of IR-MCL in simulations, IR-MCL exhibits inconsistent performance in terms of ATE and requires retraining for each new dataset. Similarly, in real environments (RE1-RE4), our method maintains the highest convergence rate among all baselines, leveraging its ability to detect deviated elements in the environment. The only exception is in RE5, where our algorithm fails due to insufficient room detection for effective graph matching, because of noisy sensor data. Moreover, AMCL and UKFL’s heavy reliance on accurate initial position estimates leads to degraded convergence rates in environments with complex geometries. OGM2PGM’s accuracy suffers particularly in symmetric environments. Our method overcomes these limitations through its robust deviation detection capabilities, though it requires a sufficient number of distinguishable rooms to function effectively.
@@
Convergence and Computation Time. Table IV shows the convergence time for each algorithm. The convergence time is the time it takes the algorithm to globally localize. IR-MCL and OGM2PGM have the best convergence time. iS-Graphs and diS-Graphs have considerably longer convergence times because they need to detect a certain number of rooms in the environment for graph matching to find a unique match. In addition, the previous version of graph matching struggled to resolve symmetries during the matching process, resulting in longer convergence times. However, the modifications in the graph matching algorithm (Section IV) proposed in this work improve the symmetry resolution ability and reduce the convergence time by almost .
Table IV shows the computation time for each algorithm. The computation time is the time used for each pose update. On average, the computation time of our algorithm is milliseconds, showing its real-time performance. Our computation time is the best compared to other 3D algorithms. OGM2PGM and AMCL have the best computation time because, unlike others, they process 2D information.
@@
Deviation Estimation. Fig. 8 shows the amount of deviation our algorithm can correctly estimate in real environments. The minimum detectable deviation is bounded by our LiDAR sensor’s accuracy. With our LiDAR having an accuracy of [27] and considering that of measurements fall within three standard deviations () of the mean, our system can reliably detect deviations only when they exceed . The maximum translational and rotational deviation in wall-surfaces our algorithm can detect accurately is and respectively. Fig. 5 shows several qualitative results from our diS-Graphs method for real experiments. The robot can successfully localize, map, and estimate deviations in these environments. Note that we only show the rooms and walls for better understanding, and all the other semantic elements and the robot are not shown. We do not show RE5 in Fig. 5, as the robot could not localize itself in this sequence.
Robustness against Deviations. Figure 7 shows the effect of deviations on the localization performance of iS-Graphs and diS-Graphs. The results consistently show that diS-Graphs, which incorporates deviation detection and modeling, maintains a relatively stable ATE even as the deviation increases from cm to cm. In contrast, iS-Graphs, which lacks explicit deviation modeling capabilities, exhibits a clear upward trend in ATE as deviations increase.
@@
Ablation Study. Table I shows that associating ‘uniform covariance’ in simulated datasets (diS-Graphs (UC)) the algorithm cannot differentiate between deviated and non-deviated elements which results in poor pose estimation. Fig. 7 demonstrates that iS-Graphs with uniform covariance shows higher error variability compared to our method which uses deviation probability-based covariance assignment. In some cases, the use of uniform covariances results in even worse performance than iS-Graphs. Similarly, when using single optimization (diS-Graphs (SO)) instead of alternating optimization, the algorithm cannot differentiate between the transformation between two graphs and the deviations between their elements, resulting in higher ATE as shown in Table I. Table II shows the ablation of uniform covariances and single optimization in real-world datasets. Using uniform covariances (diS-Graphs (UC)) and single optimization (diS-Graphs (SO)) results in lower mapping accuracy due to the algorithm’s inability to accurately map the deviated elements and differentiate between initial transformation and deviations simultaneously.
@@ VIII Conclusions and Future Work
This paper presents a novel approach to tightly couple SLAM with imprecise architectural plans. Our algorithm establishes direct correspondences between structural elements in “as-built” environments and their “as-planned” counterparts, enabling simultaneous robot localization and deviation estimation between the two representations. By detecting and estimating the deviations between “as-built” and “as-planned” environments through tight coupling, our method outperforms the current best approach with 43% better localization in simulations, 7% improved mapping accuracy in real environments, and enhanced robustness to architectural deviations. Additionally, our algorithm provides an estimate of existing deviations up to in translation and in rotation. Our algorithm is limited by the need to have enough distinctive semantic elements (i.e. wall-surfaces and rooms) to provide a unique match. As future work, we plan to add the ability to detect and match more semantic elements, which will translate into an improvement in the convergence rate and deviation detection range. Moreover, to gain flexibility, we plan to improve the matching process by adding the ability to detect rooms consisting of more than four walls.